OpenStudy (anonymous):

Find a polynomial f(x) of degree 3 that has the following zeros: -9,2,0 Leave your answer in factored form.

OpenStudy (mathmate):

If a 3rd degree polynomial has three known roots x1,x2,x3, then the factored form is: a(x-x1)(x-x2)(x-x3)=0 where a is an arbitrary constant.

OpenStudy (anonymous):

So in this case, x(x+9)(x-2) ?

OpenStudy (phi):

remember, zeros mean when x= -9 , x= 2 and x=0 then the polynomial is 0 if you rearrange these 3 equations you get x+9=0 x-2=0 x=0 when you multiply them together you get x(x-2)(x+9) = (obviously) 0 you could also multiply by some number a, and it will still be 0 a x (x-2) (x+9) = 0

OpenStudy (phi):

for this question, leave a=1 for the simplest form. But you could pick any a you want (except a=0)

OpenStudy (anonymous):

Thanks for the help.